OpenStudy (anonymous):

the length of a rectangle is 9/49 square foot times 1/9 which is the width is what?

OpenStudy (anonymous):

do u mean the area = 9/49 sq ft??

OpenStudy (anonymous):

yea i need to find the area

OpenStudy (anonymous):

ok 9/49 is length and 1/9 is width so area = 9/49 * 1/9 = 9 / 441 sq ft = 1 /49 sq ft

OpenStudy (anonymous):

so its 1/49?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

area = length times width

OpenStudy (anonymous):

oh ok much appericated

OpenStudy (anonymous):

plz click on good answer to close the question